5. use NPV and internal rate of return criteria to advise the company on whether the machine should be acquired or rejected.
b) Thelma holds two bonds; a 30year coupon bond with a face value of $30million. Suppose interest rate increased from 8% to 9%, which bond will have the greater decline in price?
1
Expert's answer
2020-11-17T16:51:53-0500

BP = "\frac{C\times(1-(1+r)^{-n}}{i} + M(1+r)^{-n}"

C = F/n = 30/30 = $1million


i. When r = 8%

BP = "\frac{1(1-(1+8\%)^{-30}}{0.08} + 30(1+0.08)^{-30}"

BP = 11.25778 + 2.98132 = $ 14.2391


ii. When r=9%

BP = "\frac{1(1-(1+9\%)^{-30})}{0.09} + 30(1+0.09)^{-30}"

BP = 10.27365 + 2.261134 = $ 12.53479

The bond with a great decline in price is the bond with an interest rate of 9% (bond price = $ 12.53479).